### v2

# Verbosity-Aware Rationale Reduction: Effective Reduction of Redundant Rationale via Principled Criteria

Abstract
Large Language Models (LLMs) rely on generating extensive intermediate reasoning units (e.g., tokens, sentences) to enhance final answer quality across a wide range of complex tasks. While generating multiple reasoning paths or iteratively refining rationales proves effective for improving performance, these approaches inevitably result in significantly higher inference costs. In this work, we propose a novel sentence-level rationale reduction training framework that leverages likelihood-based criteria, verbosity, to identify and remove redundant reasoning sentences. Unlike previous approaches that utilize token-level reduction, our sentence-level reduction framework maintains model performance while reducing generation length. This preserves the original reasoning abilities of LLMs and achieves an average 17.15% reduction in generation costs across various models and tasks.
Verbosity-Aware Rationale Reduction:
Effective Reduction of Redundant Rationale via Principled Criteria
Joonwon Jang1  Jaehee Kim 2  Wonbin Kweon3  Hwanjo Yu1,4,∗ POSTECH1,3,4, Seoul National University2 {kaoara, skwb4453, hwanjoyu}@postech.ac.kr 1,3,4 jaehee_kim@snu.ac.kr 2
1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities comparable to human cognitive abilities (Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). These models can solve complex reasoning tasks by explicitly generating extended reasoning paths. The generation of such paths involves producing explicit reasoning units (e.g., tokens, steps) (Yu et al., 2024b), which additionally enhances model performance through iterative prompting (Wang et al., 2023; Yao et al., 2023). This iterative generation of explicit reasoning paths allows the model to refine and expand its thought processes, incorporating strategic planning and ongoing cognitive generation (Xi et al., 2023; Yang et al., 2024).
Although the extensive generation of explicit reasoning units leads to better peformance, it inherently leads to higher inference costs and increased latency (Yu et al., 2024b; Wang et al., 2024). Moreover, fine-tuning LLMs using a complete reasoning path does not consistently guarantee stronger performance (Yu et al., 2024b; Deng et al., 2024b; Liu et al., 2024), suggesting the need for methods that can maintain reasoning performance while reducing the generation of reasoning units. Despite this apparent requirement, it remains underexplored how to maintain LLM reasoning capabilities while reducing intermediate reasoning paths across diverse tasks.
Previous methods (Yu et al., 2024b) have drawn inspiration from human cognitive processes to address the aforementioned limitations. Some studies have proposed training pipelines that utilize augmented datasets, iteratively generated by foundation LLMs, to fine-tune subsequent LLMs (Yu et al., 2024b; Liu et al., 2024). However, they are inherently vulnerable as they significantly depend on the generative capabilities of LLMs. In response, other works have directly trained LLMs without augmenting datasets that facilitate implicit reasoning before answer generation. Deng et al. (2023) introduced a knowledge distillation method to distill explicit reasoning into implicit reasoning through token-level hidden states. Deng et al. (2024b) selected tokens as the reasoning unit for reduction and proposed a heuristic method to internalize explicit intermediate rationale tokens. They progressively eliminate tokens from the beginning of the full rationale during the Chain of Thought (CoT) fine-tuning. However, these methods present a fundamental limitation as they lack sufficient justification for selecting tokens over more linguistically natural units (e.g., sentences) for reduction (Table 1), and they fail to provide principled criteria for the removal process. Furthermore, their evaluation has been predominantly conducted on synthetic arithmetic reasoning tasks, which presents limitations for real-world applications.
To address these limitations, we propose a novel training method for maintaining LLM reasoning performance while systematically eliminating redundant reasoning units within the reasoning process. Our method leverages sentences as fundamental reduction units, establishing more linguistically coherent boundaries compared to token-level approaches. We first empirically demonstrate that rationale sentences in early reasoning stages can introduce redundancy in LLM’s answer generation process. Drawing inspiration from Dong et al. (2023), we introduce the concept of ‘verbosity’ and develop a CoT fine-tuning method that removes unnecessary rationale sentences, thereby reducing rationale generation while preserving reasoning capabilities.
The primary contributions of this work are as follows:
  1. 1.
    We propose a novel method for sentence-level reasoning path reduction in LLMs.
  2. 2.
    We demonstrate that principled criteria can effectively identify redundancies and reduce intermediate token generation efficiently.
  3. 3.
    We validate our method’s effectiveness and robustness through extensive experiments across various LLMs and ablation studies.
2 Related Works
2.1 Performance-Cost Tradeoffs in Reasoning Path Generation
Recent research has demonstrated the importance of generating iterative and refined reasoning paths to improve the reasoning capabilities of models, albeit at increased computational costs (Wang et al., 2023; Yao et al., 2023; Radha et al., 2024; Wang et al., 2024; Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). Self-Consistency (Wang et al., 2023) improves reasoning accuracy by generating more than 20 distinct reasoning paths and selecting the most consistent answer using ensemble methods, resulting in substantial computational overhead. Similarly, Tree of Thoughts (ToT) (Yao et al., 2023) expands the exploration of reasoning paths using a tree-structured method, while Strategic Chain of Thought (SCoT) (Wang et al., 2024) utilizes an Inner Dialogue Agent to adaptively explore the diverse reasoning path.
Concurrently, Self-Refine (Madaan et al., 2024) utilizes an iterative framework that incorporates feedback to refine responses and enhance reasoning accuracy, necessitating multiple forward passes through the model. Additionally, the Reflexion-based framework (Shinn et al., 2024; Kumar et al., 2024) guides models to iteratively generate relevant reasoning paths through reflection for accurate answer generation. Given the necessity of iterative generation and refinement of reasoning paths to achieve optimal performance, they inherently increase inference costs and latency. Therefore, it is essential to investigate methods for efficiently generating these paths.
2.2 Reasoning Path Reduction
To address the computational costs associated with extensive reasoning paths generation, some lines of work (Yu et al., 2024b; Liu et al., 2024) have focused on generating augmented datasets with varying rationale lengths to shorten the generation of reasoning paths. Yu et al. (2024b) guide model to generate multiple reasoning paths using Self-Consistency to construct datasets, subsequently fine-tuning the model to generate direct answers without intermediate reasoning steps. Liu et al. (2024) developed a heuristic approach to merge reasoning steps and iteratively trained the model to produce shorter reasoning paths, which are subsequently incorporated into the progressive training phase. While these approaches demonstrate empirical effectiveness, they present two fundamental limitations: (1) their heavy dependence on LLM generation capabilities introduces inherent instability, and (2) their objective of reducing reasoning path necessitates the paradoxical creation of datasets requiring extensive reasoning path generation.
To address these weaknesses, another line of work (Deng et al., 2023, 2024b) has focused on directly training LLM without augmented datasets. Implicit-CoT (Deng et al., 2023) utilizes a multi-model framework where an emulator model is trained to predict the teacher’s token-level hidden states, and a student model uses these predicted states to generate answers. ICoT-SI (Deng et al., 2024b) identifies tokens as reduction units, proposing a method to internalize explicit intermediate rationale tokens by progressively removing them from the beginning of the reasoning path within the CoT fine-tuning process. However, these methods have shown limited generalization across diverse datasets as they have only been validated on simple arithmetic reasoning tasks, such as multiplication problems. This limitation raises concerns about their applicability to real-world scenarios where rationales are expressed in natural language. Moreover, they do not justify their criteria for token-level removal nor consider whether more linguistically natural units (e.g., sentences) might serve as more effective reduction units. Specifically, the token-level removal approach may inadvertently eliminate critical information necessary for answer generation or distort the semantic information of the sentence, while requiring repeated removal of redundant reasoning steps. Motivated by the computational costs of the extensive reasoning path generation and the limitations of existing reduction approaches, we investigate the redundancy of various sentence positions for potential removal and propose a novel method with principled criteria that can efficiently reduce reasoning paths while maintaining their effectiveness.
3 Early Stage Rationales are Redundant
Figure 1: NLL differences across varying sizes of $\{r_{i}\}$. The ‘original’ represents the NLL for the full rationale, while ‘front’, ‘random’, and ‘back’ indicate that $\{r_{i}\}$ is sampled from the front, random, and back indices of the full index set, respectively.
3.1 Quanitifying the redundancy
Before delving into the method, we first investigate which positions within the rationale sentences should be selected for reduction. When the likelihood of the answer remains unchanged after a sentence is removed from the full rationales, it indicates that the sentence may be redundant for the reasoning process. To quantify the redundancy of a sentence, we compute the negative log-likelihood (NLL) for answer $y$ after the sentence reduction as follows:
$\begin{split}\text{NLL}&=-\log p_{\theta}(y|R^{\prime},x),\\ \text{where}\ R^{\prime}&=R\setminus\{r_{i}\}_{i\in S},\ S\subseteq I.\end{split}$ (1)
Let $R$ denote the complete set of rationale sentences, and $I$ represent the full index set of these sentences. The subset of indices corresponding to sentences selected for reduction is denoted by $S$, and $R^{\prime}$ represents the remaining rationale sentences after their removal. $\{r_{i}\}_{i\in S}$ denotes the sentences selected for reduction. For simplicity, we use $\{r_{i}\}$ without the subset index notation throughout the rest of the paper.
3.2 Redundancy of leading sentences
We conduct a pilot study to empirically demonstrate the redundancy of the leading sentences within the rationales. By varying the size of $\{r_{i}\}$ from 1 to 4, we calculate the NLL of Mistral 7B (Jiang et al., 2023) across different reasoning datasets. To explore more general patterns of sentence removal, we employ a stochastic strategy in our experimental setup. We compare three configurations for sampling $\{r_{i}\}$: front, random, and back. In the front configuration, initial sentences are assigned a higher probability when forming $R^{\prime}$, whereas random and back configurations use uniform and progressively increasing probabilities, respectively [^1] For generating $R^{\prime}$, we assign probabilities $p_{k}=\frac{N-k+1}{\sum_{i=1}^{N}i}$ (front), $\frac{1}{N}$ (random), and $\frac{k}{\sum_{i=1}^{N}i}$ (back) where k=1,…,N denotes sentence position.. Additionally, we computed the NLL for full rationale sentences (i.e., $-\log p_{\theta}(y|R,x)$) as a baseline to evaluate the impact of reduction. As depicted in Figure 1, the NLL difference between the full and reduced rationales shows a marginal likelihood difference in the front ($<$) configuration. In contrast, removing sentences randomly ($\diamond$) or from the back ($>$) results in higher NLL as the removed sentences increase, highlighting the importance of the selection of removal position in the reasoning and answer prediction process (for additional data, see Appendix A).
4 Verbosity-Aware Rationale Reduction
Figure 2: Overview of our VARR/VARR+ framework. Initially, we select a candidate sentence from the beginning of the rationale (Section 3). After removing the candidate sentence, we evaluate $\textit{verbosity}(y_{g})$ and $\textit{verbosity}(y_{w})$ using Equations (5) and (9). If the candidate sentence meets the verbosity evaluation criteria, it is excluded in subsequent training steps. The model then proceeds with training, where the redundant sentence is excluded from the rationale.
Based on the above observation, we propose the Verbosity-Aware Rationale Reduction (VARR) framework. In Section 4.1, we introduce the concept of ‘verbosity’ as a principled criterion for identifying redundant reasoning sentences. Section 4.2 describes the integration of verbosity in the reduction process during CoT training. Subsequently, Section 4.3 presents additional verbosity term by incorporating the wrong answer to improve robustness. Finally, Section 4.4 illustrates the overall VARR framework.
4.1 Verbosity
To quantify the redundancy of the sentences for removal, we first introduce a foundational concept ‘verbosity’. Given an input $x$, full rationale $R$, and a reduced rationale $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, we measure the verbosity of a sentence $r_{i}$ on $y$ by measuring the difference in Kullback–Leibler divergence (KL-divergence) as follows:
$\begin{split}\textit{verbosity}(y)&=D_{KL}(q(y|x)\parallel p_{\theta}(y|R,x))\\ &\hskip 15.00002pt-D_{KL}(q(y|x)\parallel p_{\theta}(y|R^{\prime},x)),\end{split}$ (2)
where $q(y|x)$ is the ground truth distribution. The verbosity$(y)$ highlights the informational contribution or redundancy of a rationale sentence $r_{i}$ on answer $y$. Since $q(y|x)$ is the form of the one-hot vector (i.e., Dirac delta function), the verbosity$(y)$ can be computed as the log-likelihood ratio between $R$ and $R^{\prime}$ as follows:
$\begin{split}\textit{verbosity}(y_{g})&=\left[H_{q}(p_{\theta}(y|R,x))-H(q(y|x))\right]\\ &\hskip 15.00002pt-\left[H_{q}(p_{\theta}(y|R^{\prime},x))-H(q(y|x))\right]\\ &=E_{q}[-\log P(y|R,x)]\\ &\hskip 15.00002pt+E_{q}[\log P(y|R^{\prime},x)]\\ &=\log\left(\frac{P(y_{g}|R^{\prime},x)}{P(y_{g}|R,x)}\right),\end{split}$ (3)
where $y_{g}$ is the ground truth answer (i.e., $q(y_{g}|x)=1$). $H_{q}(\cdot)$ and $H(\cdot)$ denote the cross-entropy and the entropy, respectively. Intuitvely, a higher value of $\textit{verbosity}(y_{g})$ implies that the likelihood of the model generating the ground truth answer increases after removing $r_{i}$, indicating that its removal is beneficial.
4.2 Verbosity Identification in CoT Training
Given an input sequence, the CoT training (Nye et al., 2021) aims to train LLMs to generate full rationale, followed by the ground truth answer with the following loss:
$-\log p_{\theta}(y_{g},R|x).$ (4)
In each training step $t$, each sentence $r_{i}$ within $R$ is assessed with the following the equation:
$\textit{verbosity}(y_{g})\geq 0.$ (5)
Here, we sequentially select $r_{i}$ starting from the first sentence and construct $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, based on our analysis in Section 3, which indicates that early-stage rationales are more likely to be redundant. A non-positive verbosity score indicates that removing $r_{i}$ from $R$ impairs the model’s performance, thereby identifying the sentence as essential and retaining it in the rationale.
4.3 Contrasting with Wrong Answer
Inspired by the miscalibrated log-likelihoods between accepted and rejected responses during standalone Supervised Fine-Tuning (SFT) in alignment learning Rafailov et al. (2024); Azar et al. (2024); Hong et al. (2024), we also assess whether the reduced rationales $R^{\prime}$ result in inaccuracy answer generation by introducing a wrong answer $y_{w}$. Instead of employing the correct answer distribution $q(y|x)$ in Equation (2), we initiate our formula with the wrong answer distribution $q^{\prime}$ (i.e., $1-q(y|x)$, normalized sum to 1) Through algebraic manipulation, we can obtain:
$\begin{split}&D_{KL}(q^{\prime}(y|x)\parallel p_{\theta}(y|R,x))\\ &\hskip 10.00002pt-D_{KL}(q^{\prime}(y|x)\parallel p_{\theta}(y|R^{\prime},x))\\ &=\left[H_{w}(p_{\theta}(y|R,x))-H(q^{\prime}(y|x))\right]\\ &\hskip 10.00002pt-\left[H_{w}(p_{\theta}(y|R^{\prime},x))-H(q^{\prime}(y|x))\right],\end{split}$ (6)
where $H_{w}(\cdot)$ denotes the cross-entropy calculated with the wrong answer distribution $q^{\prime}$. Due to the impracticality of computing the expectation over the entire space of $V-1$ wrong answers (where $V$ is the vocabulary size), we sample $K$ incorrect answers for the following estimations:
$\begin{split}&E_{w}[-\log p_{\theta}(y|R,x)]+E_{w}[\log p_{\theta}(y|R^{\prime},x)]\\ &={E}_{w}\Big{[}\log\frac{p_{\theta}(y|R^{\prime},x)}{p_{\theta}(y|R,x)}\Big{]}\\ &\approx\frac{1}{K}\sum_{k=1}^{K}\log\frac{p_{\theta}(y_{w}^{(k)}|R^{\prime},x)}{p_{\theta}(y_{w}^{(k)}|R,x)},\end{split}$ (7)
where $\{y_{w}^{(k)}\}_{k\in\left[K\right]}$ is sampled from the in-batch negatives depending on the dataset. Consequently, the $\textit{verbosity}(y_{w})$ is computed as:
$\textit{verbosity}(y_{w})=\frac{1}{K}\sum_{k=1}^{K}\log\frac{p_{\theta}(y_{w}^{(k)}|R^{\prime},x)}{p_{\theta}(y_{w}^{(k)}|R,x)}.$ (8)
Since computational constraints necessitate sampling incorrect answers to calculate $\textit{verbosity}(y_{w})$, we contrast $\textit{verbosity}(y_{w})$ with $\textit{verbosity}(y_{g})$ rather than solely using $\textit{verbosity}(y_{w})$. This ensures that the likelihood of incorrect answers remains lower than that of ground truth answers, even after the removal of $r_{i}$ as follows:
$\textit{verbosity}(y_{w})-\textit{verbosity}(y_{g})\leq 0.$ (9)
A smaller $\textit{verbosity}(y_{w})$ value compared to $\textit{verbosity}(y_{g})$ indicates that $R^{\prime}$ increases the likelihood of generating incorrect answers over the ground truth answer after removing $r_{i}$, making it beneficial to remove.
4.4 CoT Training with Rationale Reduction
In our framework, the model is trained using Equation (4) for predefined warm-up steps to inject its reasoning capabilities. After the warm-up steps, at each training step $t$, each sentence $r_{i}$, starting from the first index sentence in $R$, is evaluated against either Equation (5) alone (denoted as VARR) or both Equations (5) and (9) (denoted as VARR+). Sentences satisfying these respective criteria are selected for removal and are not included in subsequent training steps. The maximum removable number of sentences at each training step $t$ is determined based on a linear schedule, adopting Deng et al. (2024b)’s setting as follows:
$r(t)=\lfloor N_{t}\cdot(t/T)\rfloor,$ (10)
where $T$ represents the total number of training steps, $N_{t}$ is the total number of rationale sentences at step $t$, and $r(t)$ indicates the maximum number of sentences that can be removed at that step.
5 Experiments
5.1 Training Configuration
Figure 3: Performance for 5 different reasoning tasks. Acc refers to Accuracy while Tokens denotes the Generated Tokens. Bar plots illustrate the accuracy of each method on datasets, while line plots show the average generated tokens by each method across the datasets. The y-axis represents the range of each metric, and the x-axis lists the datasets evaluated.
Datasets
We conducted experiments across two categories to provide a comprehensive evaluation of VARR’s versatility and effectiveness, in contrast to prior research that predominantly focuses on simple arithmetic tasks like multi-digit multiplication (Deng et al., 2023, 2024b). Initially, we evaluate with arithmetic reasoning tasks, including datasets like GSM8K (G8K; Cobbe et al. 2021a) and MathQA (MQA; Amini et al. 2019a). We also examine the performance of our method on commonsense reasoning tasks, employing datasets including CommonsenseQA (CQA; Talmor et al. 2019), TrivaiQA (TQA; Joshi et al. 2017), and StrategyQA (SQA; Geva et al. 2021).
Models
We trained Mistral 7B (Jiang et al., 2023) as our base model for comparisons. We also trained a series of Llama3.2 models (AI@Meta, 2024) scaling from 1B to 3B to demonstrate our method’s generalization capabilities.
Implementation Details
The warm-up stage is set to 0.1 of the total training steps, and its impact is analyzed in Section 5.5.4. In addition, an optimizer is reinitialized at the beginning of each epoch to stabilize the model training inspired by Deng et al. (2024b)’s setting, with its effects described in Section 5.5.3. All baselines and our proposed method are trained for 5 epochs to ensure fairness in comparison. Detailed implementation of VARR/VARR+ can be found in the Appendix B.
5.2 Baselines
We compare our method to the following baselines: Explicit-CoT (Nye et al., 2021), where the model is finetuned with explicit chain-of-thought reasoning, and ICoT-SI (Deng et al., 2024b), where the model is fine-tuned using a linear token elimination schedule. It is important to note that after removing intermediate reasoning paths, the model’s performance should be competitive with Explicit-CoT. Consequently, ICoT-SI serves as our primary baseline. We exclude Implicit-CoT Deng et al. (2023) from our baselines due to its significant memory requirements, requiring the training of three models on a single GPU. Furthermore, according to Deng et al. (2024b), this method underperformed compared to ICoT-SI. All baselines were trained on a single A100-80GB GPU, and detailed training configurations for each method are provided in Appendix C.
5.3 Evaluation
We employ two evaluation metrics: First, we evaluate the accuracy of each method in generating the final answer for the respective tasks. Second, we count the generated tokens to evaluate reasoning efficiency while maintaining performance.
5.4 Main results
In Figure 3, we present the results for each reasoning task across different models. Our method demonstrates consistent performance improvements or maintains comparable results across most benchmarks, while also achieving more efficient reasoning path generation. Notably, we observed significant improvements in average accuracy with increments of 2.59%, 6.27%, and 14.27% for Mistral 7B, Llama3.2 1B, and Llama3.2 3B (VARR+), respectively, compared to Explicit CoT, while efficiency improved by 19.35%, 17.24%, and 14.88% for these models. These findings indicate that some reasoning sentences are redundant and can even disturb answer generation, highlighting that effective reasoning capabilities can be learned through selective utilization of essential reasoning units.
For ICoT-SI, the model exhibited performance degradation after token removal training. Specifically, ICoT-SI shows an average performance degradation of 18.93% compared to Explicit CoT. Furthermore, while the Mistral 7B model achieved a 50.26% accuracy on the GSM8K using a large augmented dataset with succinctly generated reasoning paths (trained with the 385K-size dataset used in the original paper (Deng et al., 2024b)), its performance dropped significantly to 28.27% when trained on the standard GSM8K dataset (trained with the 7K-size dataset). This indicates that while ICoT-SI may excessively reduce the length of rationales in reasoning tasks, identifying the appropriate reasoning unit and applying principled criteria are crucial for maintaining performance in real-world application problems. Qualitative results across baselines are presented in Appendix F, and a detailed analysis of the reasoning unit will be discussed in Section 5.5.1.
Moreover, introducing a wrong answer in VARR led to performance improvements in most datasets. While solely using the verbosity($y_{g}$) enhanced efficient generation, it is notable that verbosity($y_{w}$) constraints reduced rationales to increase the likelihood of generating incorrect answers, which contributes to the robustness and stability of the training process.
5.5 Ablation Studies
In this section, we conduct ablation studies to empirically validate our method. All experiments are implemented using Mistral 7B due to its higher base capacity compared to other models.
5.5.1 Identifying Appropriate Units for Removal
G8K MQA CQA TQA SQA
Exp-CoT 55.26 55.84 84.33 82.94 74.70
(122.54) (105.02) (42.84) (63.46) (46.47)
ICoT-SI 28.27 35.84 67.82 77.09 61.33
(104.41) (113.55) (6.0) (22.03) (22.86)
VARR-Tok 47.53 46.79 82.60 67.14 71.22
(94.16) (90.39) (25.07) (38.10) (31.92)
VARR-Sent 54.98 56.95 89.56 83.45 78.19
(100.38) (91.04) (33.55) (39.17) (35.12)
Table 1: Analysis across various reasoning reduction units and the application of principled criteria. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
In this section, we experimentally demonstrate why the sentences are suitable units for reduction reasoning units. Unlike ICoT-SI which lacks specific criteria for token removal, we apply VARR+ at the token level (denoted as VARR-Tok) to assess whether tokens can effectively function as reduction units in reasoning processes. As depicted in Table 1, applying principled criteria at the token level results in an average performance gain of 26.95%, indicating that the principled criteria contribute to robust performance. Moreover, extending the reasoning units from tokens to sentences (denoted as VARR-Sent) leads to a further performance increase of 21.83% over VARR-Tok, highlighting the importance of selecting the appropriate reasoning unit for effective reasoning path reduction. These findings demonstrate the importance of selecting appropriate reasoning units, with sentences serving as linguistically natural and effective boundaries for the removal process.
5.5.2 Analyzing the Impact of Sentence Position on Removal Efficacy
G8K MQA CQA TQA SQA
Exp-CoT 55.26 55.84 84.33 82.94 74.70
(122.54) (105.02) (42.84) (63.46) (46.47)
No Rule 25.85 35.06 74.78 74.14 72.38
(63.56) (59.92) (23.56) (6.84) (23.38)
Random 52.31 55.34 83.47 79.62 73.83
(115.59) (99.13) (49.19) (44.98) (50.51)
Back 48.52 49.71 85.21 70.0 74.41
(103.25) (92.64) (19.86) (36.22) (18.17)
Front 54.98 56.95 89.56 83.45 78.19
(100.38) (91.04) (33.55) (39.17) (35.12)
Table 2: Performance across 5 different reasoning tasks, evaluated with different sentence position removal. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
In Section 3, we demonstrated that gradually removing sentences from random and back positions can degrade model performance. To further explore this finding and assess the robustness of removing sentences from the front position, we conducted experiments with unguided random sentence removal (denoted as No Rule) and applied VARR+ with random position and reverse sentence order (denoted as Random and Back, respectively). As shown in Table 2, unguided random sentence removal resulted in a 25.30% decrease in performance relative to our method, highlighting the critical role of systematic configuration even after selecting sentences as the units of reduction. Additionally, Random and Back strategies result in an average 7.50% performance degradation compared to our method. This finding again indicates that earlier sentences in the reasoning path tend to be more redundant, and prioritizing their removal effectively balances rationale reduction while maintaining reasoning performance.
5.5.3 Reinitializing the Optimizer
G8K MQA CQA TQA SQA
w/o reinit 45.18 48.70 87.82 83.72 76.74
(115.91) (92.86) (32.95) (38.27) (33.89)
w/ reinit 54.98 56.95 89.56 83.45 78.19
(100.38) (91.04) (33.55) (39.17) (35.12)
Table 3: Performance Comparison before and after the reinitializing optimizer in every training epoch. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
We reinitialized the optimizer after each training epoch to stabilize training, inspired by Deng et al. (2024b). We employed the AdamW Loshchilov and Hutter (2019) optimizer, where the first and second moments are gradually updated based on current gradients, which could lead to unstable training in our framework. Specifically, after applying VARR throughout the dataset, a reduction in the rationale for some data between epochs could lead to instability in the training process. With a 19.32% performance gain achieved by reinitializing the optimizer, we reconfirm that Deng et al. (2024b)’s findings are robust across diverse datasets with complex semantic and syntactic structures in reasoning paths.
5.5.4 Varying the Warm-up Steps
Figure 4: Accuracy and Generated Token difference across various warm-up stages.
We evaluated our method against various warm-up stages on the TriviaQA and CommonsenseQA datasets. As the duration of the warm-up stages increases, the model becomes more fitted to the non-reduction dataset, which inhibits VARR’s ability to eliminate redundant sentences from the reasoning path. Consequently, as illustrated in Figure 4, longer warm-up periods result in an increase in generated tokens and a decrease in accuracy. These results suggest that 0.1 training steps provide sufficient time to inject reasoning abilities while enabling the systematic reduction of redundant reasoning sentences during the learning process. Therefore, we set 0.1 as the default setting for the warm-up steps.
5.5.5 Removal Ratio Analysis
Figure 5: Removal ratio of redundant sentences throughout the training stage. The Removal Ratio, indicated on the y-axis, is calculated by dividing the number of sentences actually removed by the maximum potential removal ratio (i.e., ${size(\{r_{i}\})/r(t)})$. The x-axis represents the index of the epochs.
In Figure 5, we analyze the actual amount of rationale sentences removed during training. We examine it by calculating the removal ratio, the proportion of actual removed sentences to the maximum potential removal sentences $r(t)$. Our analysis indicates that not all sentences designated for maximum removal range are always eliminated during the training process. Notably, a significant proportion of redundant sentences are removed in the early stages of training, with fewer sentences being removed as the model progresses through the middle to later stages, thereby stabilizing its training. A similar trend is observed across other datasets, as detailed in Appendix E.
6 Conclusion
In this work, we propose sentence-level rationale reduction framework VARR and provide empirical evidence that models trained with non-redundant rationales achieve greater efficiency. We address the lack of principled criteria for identifying redundant sentences during the training phase by developing a reduction method that not only preserves the model’s reasoning capabilities but also reduces the likelihood of generating incorrect answers. Our experiments show that VARR can efficiently handle a diverse range of tasks with fewer generated tokens, without sacrificing its accuracy. Further empirical findings confirm that sentences are appropriate reasoning units for rationale reduction, highlighting the necessity of principled criteria for robust application. This work contributes novel insights to rationale reduction research, contributing to the efficient elicitation of reasoning in language models.
Limitations
While our work provides novel insights into rationale reduction research, our experiments were primarily conducted using a relatively small large language model, constrained by computational costs. Additionally, for the same reasons, it was not feasible to test the model with datasets featuring long sequences in both queries and rationales (Reddy et al., 2024; Yu et al., 2024a). However, given that the VARR/VARR+ frameworks are designed with a systematically structured approach, we believe that they could also be effective in scaled-up settings.
Appendix A Additional Datasets Analysis
Figure 6: NLL differences across varying sizes of $\{r_{i}\}$. The ‘original’ represents the NLL for the full rationale, while ‘front’, ‘random’, and ‘back’ indicate that ${r_{i}}$ are sampled from the front, random, and back indices of the full index set, respectively.
As shown in Section 3, the Negative Log-Likelihood (NLL) tends to rise with increasing size of ${r_{i}}$ in random and back configurations, compared to the front configuration, as illustrated in Figure 7 this trend is consistent across both the GSM8K and StrategyQA datasets.
Appendix B VARR Implementation
The detailed implementation of VARR+ is outlined in Algorithm 1. After the warm-up stages, each datum in the current batch is evaluated using the verbosity equation—specifically, using only Equation 5 in VARR and both Equations 5 and 9 in VARR+. As mentioned in Section 3 and 4, the process assesses the redundancy of each sentence in data starting from the first index during every single epoch. Therefore, all data can potentially shorten the length of the rationale progressively.
Algorithm 1 Training Procedure of VARR+
1:$\mathcal{D}$: Training dataset
$\mathcal{B}$: Training batch
3:$E$: Total training epochs
4:$S$: Total training steps per epoch
5:$T=E\times S$: Total training steps
6:$T_{warmup}=0.1\times T$: Warm-up steps
7:$R_{B}$: A buffer to store removed sentences
8:$r(t)$: Maximum number of removable sentences at $t$
9:$N_{t}$: Number of rationale sentences at $t$
10:$\theta$: Trained model parameters
11:for epoch = 1 to E do
12:     for step = 1 to S do
13:         Sample training batch $\mathcal{B}$ from $\mathcal{D}$
14:         $t=(epoch-1)\times S+step$
15:         if $t\leq T_{warmup}$ then
16:              pass
17:         end if
18:         for each $d\in\mathcal{B}$ do
19:              $R_{B}\leftarrow\{\}$
20:              for i = 1 to $N_{t}$ do
21:                  if $r_{i}$ satisfies Equations 5 and 9 then
22:                       remove $r_{i}$ from $d$
23:                       Add $r_{i}$ to $R_{B}$
24:                       if $|R_{B}|\geq r(t)$ then
25:                           break
26:                       end if
27:                  end if
28:              end for
29:         end for
30:         Forward pass
31:         Backward pass and update $\theta$
32:     end for
33:     Reinitialize optimizer
34:end for
Appendix C Experimental Details
For all experiments, we employ the AdamW optimizer Loshchilov and Hutter (2019), configured with a weight decay of 0.005. For the Mistral 7B model, we utilize an effective batch size of 12 with gradient accumulation set to 3, while the smaller Llama3.2 models use an effective batch size of 15. A constant learning rate of $5\times 10^{-6}$ is applied across all datasets, with bfloat16 precision. For Multiple-Choice and True/False tasks, complete sets of non-correct labels are employed to configure in-batch negatives to enhance the stability of the verbosity evaluation. To ensure a fair comparison, all baselines and methods are executed on a single A100 GPU with 80GB of memory for up to 5 epochs or 36 hours, whichever is reached first.
Appendix D Details and Statistics of Datasets
Dataset Reasoning Task Source Answer Format # TRAIN. # VALID. # TEST.
GSM8K (Cobbe et al., 2021a) Arithmetic Cobbe et al. (2021b) Number 6999 472 1319
MathQA (Amini et al., 2019a) Arithmetic Amini et al. (2019b) Multi-choice 29837 4475 2985
TriviaQA (Joshi et al., 2017) Commonsense Kim et al. (2023) Natural Language 8844 552 1659
CommonsenseQA (Talmor et al., 2019) Multi-choice Kim et al. (2023) Multi-choice 609 38 115
StrategyQA (Geva et al., 2021) Commonsense Sileo (2024) T/F 832 114 344
Table 4: Comprehensive statistics of the datasets used in our experiments are provided. GSM8K and MathQA are sourced directly from their original datasets, while the remaining datasets were obtained from Kim et al. (2023) and Sileo (2024) to access complete reasoning paths.
For our experimental analysis, we carefully selected a diverse set of five datasets used in prior works (Deng et al., 2023, 2024a; Liu et al., 2024; Yu et al., 2024b; Yin et al., 2024). To ensure explicit sentence boundaries, all datasets were processed using gpt4o-mini (OpenAI, 2024) to establish clear sentence demarcation (e.g., ‘He bikes 20*2=<<20*2=40>>40 miles each day for work So he bikes 40*5=<<40*5=200>>200 miles for work’ becomes ‘He bikes 202=<<202=40>>40 miles each day for work. So he bikes 405=<<405=200>>200 miles for work’). Table 4 comprehensively outlines each dataset, including its source and the size of the training, validation, and test samples.
Appendix E Removal Ratio Analysis on additional datasets.
Figure 7: Additional analysis of the Removal Ratio on other datasets. The Removal Ratio, indicated on the y-axis, is calculated by dividing the number of sentences actually removed by the maximum potential removal ratio (i.e., ${size(\{r_{i}\})/r(t)})$. The x-axis represents the index of the epochs.
In the GSM8K, Strategy-QA, and Math-QA datasets, similar trends are observed as in Section 5.5.5. Notably, in Strategy-QA, the Removal Ratio converges more rapidly than in other datasets. This is attributed to the dataset’s characteristics, where the sentences within the rationale predominantly list simple information rather than forming complex, interrelated structures. Consequently, redundant sentences are rapidly removed, leading to a rapid decrease in further removal activities.
Appendix F Qualitative Analysis
Table 5 illustrates the generation outputs for each method on the GSM8K test set. Explicit-CoT produces a complete reasoning path for answer generation; however, its redundancy in generating ‘needs a 3 piece place setting’ leads to hallucinations within the sentence, resulting in the incorrect final answer generation. In contrast, ICoT-SI, which removes reasoning paths at the token-level, results in a generation that lacks sufficient reasoning capabilities. Conversely, VARR+ efficiently generates key information by producing ‘\$15.00’ in a single sentence, leading to more effective and efficient generation.
Question
Avery needs to buy a 3 piece place setting (dinner & salad plate and a bowl) for her holiday dinner. She’s having 12 people over for dinner. If the dinner plates cost \$6.00 each and bowls each cost \$5.00 and the salad plates cost \$4.00, how much will she spend on place settings?
Explicit-CoT
She’s having 12 people over and needs a 3 piece place setting for each so she needs 12*3 = 36 place settings.
The dinner plates cost \$6.00 each and she needs 36 of them so that comes to 6*36 = \$216.00.
The bowl costs \$5.00 each and she needs 36 of them so that comes to 5*36 = \$180.00.
The salad plates cost \$4.00 each and she needs 36 of them so that comes to 4*36 = \$144.00.
All total, the place settings will cost \$216 for dinner plates, \$180 for bowls and \$144 for salad plates for a total of 216+180+144 = \$540.00
### 540
ICoT-SI
each so 12*6 = \$72.00 She’s buying 3 pieces of place settings (dinner plates, salad plates and bowls) for \$72.00 each so 3*72 = \$216.00
### 216
VARR+
She needs 12 place settings and each setting costs \$15.00 so she will spend 12*15 = \$180.00
### 180
Ground Truth Answer
### 180
Table 5: Qualitative Example: Generation outputs on GSM8K test set.